### v1

# Using VLM Reasoning to Constrain Task and Motion Planning

Abstract
In task and motion planning, high-level task planning is done over an abstraction of the world to enable efficient search in long-horizon robotics problems. However, the feasibility of these task-level plans relies on the downward refinability of the abstraction into continuous motion. When a domain’s refinability is poor, task-level plans that appear valid may ultimately fail during motion planning, requiring replanning and resulting in slower overall performance. Prior works mitigate this by encoding refinement issues as constraints to prune infeasible task plans. However, these approaches only add constraints upon refinement failure, expending significant search effort on infeasible branches. We propose VIZ-COAST, a method of leveraging the common-sense spatial reasoning of large pretrained Vision-Language Models to identify issues with downward refinement a priori, bypassing the need to fix these failures during planning. Experiments on two challenging TAMP domains show that our approach is able to extract plausible constraints from images and domain descriptions, drastically reducing planning times and, in some cases, eliminating downward refinement failures altogether, generalizing to a diverse range of instances from the broader domain.
I Introduction
Task and Motion Planning (TAMP) [1] offers a framework for solving long-horizon robotic problems that couples symbolic decision-making with geometric feasibility, such as clearing a cluttered table or navigating between rooms where a key must first be found to open a door. These problems are compelling not only because completing tasks requires multiple steps, but also because they demand reasoning across different levels of abstraction. The robot must determine what actions to take and how to execute them. This feature makes TAMP inherently challenging as it requires reconciling discrete task structures with high-dimensional geometric constraints.
The two levels of abstraction, task and motion, enable efficient planning in long-horizon problems. At the task level, actions are represented symbolically, such as “pick,” “place,” or “navigate,” with each action specified by preconditions and effects on a set of symbolic predicates. This abstraction allows the planner to efficiently reason about possible sequences of actions without searching in continuous geometric space. However, the core difficulty is that these abstractions often cannot express critical geometric information, violating the downward refinement property [2, 3]: a valid task plan may not necessarily correspond to a sequence of feasible motion plans. For instance, a symbolic plan may prescribe grasping an object, but that object may be obstructed by another object, preventing any collision-free grasp.
Figure 1: VIZ-COAST uses VLMs and SMT-based task planning to constrain TAMP problems. Our constraints generalize to unseen problem instances within a broader domain zero-shot. VIZ-COAST introduces minimal search overhead, enabling more efficient planning than prior state-of-the-art.
PDDLStream [4] is a TAMP framework which circumvents downward refinement failures by employing a sample-first approach, in which feasible continuous action parameterizations, such as grasp poses and trajectories, are sampled prior to planning via streams, allowing task planning to be done with geometric feasibility pre-certified. However, PDDLStream expends significant effort sampling parameterizations of action groundings which may not be used in the final plan. This can become highly inefficient, as the number of possible action groundings grows exponentially with the number of objects.
Alternatively, many TAMP methods take a plan-first approach, creating a complete task plan before attempting downward refinement. Intuitively, it is far more effective to prune the search space at the discrete task level than the continuous motion level, especially in long-horizon tasks. When refinement fails, plan-first methods translate motion failures to task-level constraints before replanning. For example, upon failing to find a feasible grasp pose due to an obstructing object, a constraint may be added to enforce that the offending object must not be in that position when attempting a grasp. IDTMP [5] implements this by formulating planning as with SATPlan [6, 7] through Z3 [8], a Satisfiability Modulo Theories (SMT) solver, incrementally adding constraint clauses upon failure. COAST [9], a plan-first successor to PDDLStream, adds auxiliary predicates and entities to directly encode constraints in the PDDL description. All these existing constraint-based approaches ultimately require expending additional search effort to encounter downward-refinement failures and discover the necessary constraints, often replanning several times before finding a plan that can be successfully ground to motion. Our goal is to avoid expending this effort by adding the necessary constraints before planning even begins.
Recently, large pretrained vision language models (VLMs) have demonstrated common-sense geometric and physical reasoning capabilities [10, 11], allowing them to generalize zero-shot to a wide range of settings. For this reason, they have been employed for various facets of TAMP, including subgoal generation [12], encoding open-world preferences [13], and correcting errors [14].
In this work, we propose Visual Insight for Z3 [8]-based Constraints and Streams (VIZ-COAST), a novel approach to TAMP which leverages VLMs to foresee potential downward-refinement failures and add constraints to avoid them before planning begins, eliminating the need to encounter these failures during planning and expend effort replanning.
As shown in Fig.˜1, we design a pipeline that guides a VLM to produce constraints that directly interface with an SMT-based task planner. In contrast to COAST, whose auxiliary PDDL constructions introduce significant search overhead, the SMT-based solver allows us to seamlessly incorporate constraints while using the original PDDL description.
Through experiments on two challenging TAMP domains, we demonstrate that VIZ-COAST significantly decreases planning time by reducing or even eliminating replanning caused by downward refinement failures. We also show that our constraints consistently generalize zero-shot to unseen problem instances within a broader domain.
II Background and Related Work
Figure 2: The VIZ-COAST architecture. Our Visual Reasoning Module takes as input an example scene, consisting of an image and the geometric state, the PDDL domain description, and an example of a syntactically compliant output. It produces a Python file encoding constraints through a high-level API to the SMT-based task planner. When presented with a new problem instance, the task planner takes these constraints as input, in addition to the PDDL domain and new problem descriptions. The planner applies the constraints to block geometrically infeasible task plans. The high-level plan produced is then grounded by a streams-based motion planner to produce an executable continuous motion plan.
In this section, we begin with a general overview of modern task and motion planning research and its main challenges in Sec.˜II-A. We then review prior approaches in Sec.˜II-B that accelerate TAMP by constraining the search space, highlighting the need for a generalizable method capable of pruning before planning begins. Lastly, in Sec.˜II-C, we explain how VLMs’ spatial reasoning capabilities make such a method possible and clarify how our approach differs from prior work.
II-A Task and Motion Planning
Task and motion planning (TAMP) decomposes continuous long-horizon motion planning into search in both discrete (task-level) and continuous (motion-level) spaces [1]. The task level planning is done over an abstraction of the robot’s capabilities, and produces a sequence of symbolic actions to transition from a given start state to a desired goal condition (e.g., STRIPS [15]). In this work we formalize task planning using Planning Domain Definition Language (PDDL) [16] due to its expressive power and widespread adoption.
A core challenge in TAMP is downward-refinement failure [2, 3], which occurs when a symbolic task plan becomes infeasible at the motion level. PDDLStream [4] addresses this issue by taking a sample-first approach, performing continuous sampling before task planning via declarative streams, allowing earlier reasoning about geometric feasibility. Other works [17, 18, 19, 9] take a plan-first approach, designing algorithms to add constraints upon encountering failures which inform task-level replanning to prevent revisiting infeasible options. In the next section, we expand on the role of constraints and other methods of reducing the search space in TAMP.
II-B Constraints in TAMP
The purpose of constraints in both IDTMP and COAST is to prune the task-level search space to avoid repeating failures, though they differ in implementation. COAST [9] is a plan-first adaptation of PDDLStream which encodes constraints by modifying the PDDL description, adding auxiliary predicates and entities to block actions under specific conditions. While this approach allows COAST to be planner-agnostic, encoding constraints directly in PDDL limits the constraints that may be expressed, and the auxiliary constructions introduce significant search overhead. PDDL 3.0 [20] introduced constraints on valid states, but is poorly supported by existing planners (e.g., Fast Downward [21]) and cannot be reliably generated by LLMs due to limited training data. In contrast, we take the approach done by IDTMP [5], and solve for valid task plans through SATPlan [6, 7] encoded in a Satisfiability Modulo Theories (SMT) solver (in particular, Z3 [8]). This formulation enables constraints to be directly encoded in the SMT solver through simple Python programming, enabling efficient and flexible constraint specification. While heuristic search approaches typically exhibit faster planning times [22], SATPlan encodings offer far greater specification flexibility and can avoid the costly translation overhead in FastDownward [21] for many problems.
Although existing methods [9, 5] are able to overcome downward refinement issues, they discover constraints only by encountering motion grounding failures during execution. This reactive process is often repeated for several iterations before uncovering the necessary constraints to solve a particular problem instance, and cannot generalize to unseen instances.
Alternative approaches have explored predicting action feasibility to avoid refinement failures. Wells et al. [23] learn feasibility constraints for tabletop manipulation through domain-specific training, while Xu et al. [24] employ neural feasibility checks to identify infeasible actions before motion planning. Driess et al. [25] use deep visual heuristics to assess feasibility in mixed-integer manipulation planning, and [26] leverage visual reasoning to predict valid actions for TAMP. Although promising, these methods require extensive domain-specific training data and may not generalize well to new environments or object configurations. In contrast, recent advances in LLMs and VLMs enable the use of pre-trained common-sense priors to reduce unnecessary exploration.
II-C Vision Language Models for TAMP
LLMs have demonstrated that they are suitable for high [27]. VLMs further integrate visual and textual inputs to enable scene understanding [28], spatial reasoning [11], and grounded action execution [29, 14]. These capabilities make VLMs especially promising for addressing the geometric reasoning challenges in TAMP.
Several works have integrated LLMs and VLMs into TAMP pipelines for various purposes. Zhang et al. [30] use VLMs to ground task planners by validating symbolic predicates and providing incremental feedback. Hu et al. [31] uses a VLM to generate physically feasible actions. ReplanVLM [32] employs VLMs for plan generation with replanning upon grounding failures. CaStL [33] translates natural language task descriptions into PDDL with constraints reflecting user preferences. OWL-TAMP [13] encodes open-world preferences as constraints on symbolic and geometric states. Both CaStL and OWL-TAMP operate at the problem specification level, aiming to capture preferences rather than prevent downward refinement failures. VLM-TAMP [12] uses VLMs’ common-sense knowledge to generate intermediate subgoals for long-horizon planning; however, it assumes plans can always be refined at the motion level, making it prone to failure when this assumption fails. CLIMB [34] incrementally updates the domain description upon failure with an LLM to continually improve task performance; however, it is restricted to constraints expressible in PDDL, and does not make use of visual input.
Our work leverages VLMs’ spatial reasoning capabilities to proactively identify general geometric constraints that prevent downward refinement failures before planning begins. This represents a novel application of VLM reasoning to the fundamental challenge of bridging symbolic and geometric reasoning in TAMP.
III Problem Formulation
Our work investigates the use of VLMs to accelerate solving the TAMP problem. In particular, our method is capable of producing constraints which generalize to a range of problem instances within a broader domain. When inferring constraints for a particular domain, we assume access to a model of the robot and its environment, as well as a Planning Domain Definition Language (PDDL) [35] domain description which specifies the available actions, predicates, and possible objects. Additionally, we assume access to an image of some example scene within the domain. We assume knowledge of the underlying geometric state of this example scene, consisting of the coordinates and type of each object present. The constraint inference step does not involve a PDDL problem description. Once constraints have been generated, they are used to plan on specific problem instances. At this stage, the specific PDDL problem description is accessed by the task planner.
We use the PDDL standard for symbolic task planning. For a given domain, a PDDL domain specifies the allowed types, predicates and actions. Actions are defined by their preconditions, expressed as logical formulas over predicates, and their effects, which specify how the truth values of a subset of predicates are updated in the next state. An instance of a problem in a particular domain defines a set of objects, their initial state, and the goal conditions. States are represented as a set of predicates which true. We adopt the closed-world assumption, in predicates not specified to be true are implicitly assumed to be false. The problem of task planning is to find a sequence of actions which achieves the goal from the initial state.
In TAMP, symbolic actions must also be grounded in the continuous domain through motion planning. For example, a symbolic action such as pick(block1, gripper) requires the motion planner to generate a feasible sequence of trajectories and determine a valid grasp pose that enables the manipulator to successfully pick up the block.
This gap between task-level validity and motion-level feasibility is the core challenge posed by TAMP that we seek to address in this work, by foreseeing and imposing constraints on task-planning to improve downward-refinability.
IV Methodology
Figure 3: VIZ-COAST’s Visual Reasoning Module infers constraints through a 4-step prompting procedure. First, it is asked to provide an interpretation of the scene based on an image of the example problem instance and the PDDL domain description. Then, it is asked to articulate in natural language the necessary constraints. After that, the VLM is asked to formally encode the constraints it identified in Python using Z3’s API, referencing a structural example. Finally, it proofreads the script to ensure syntactic compliance.
Figure 4: Paraphrased example of a constraints file produced by the Visual Reasoning Module for the Containers domain. The VRM produces a function which calls a high-level API to block action assignments under certain conditions, constraining the search space of the task planner. This code prevents any pick or place action whose target is a closed container.
VIZ-COAST consists of 3 primary components: The task planning module, the visual reasoning module (VRM), and the motion grounding module.
  • SMT-based Task planner: We use an Satisfiability Module Theories (SMT) solver—Z3 [8], in particular—that uses a SATPlan encoding [6, 7] to enable flexible constraint specification while introducing minimal search overhead. This is similar to the task planning framework used in IDTMP [5] and CaStL [33].
  • Visual Reasoning Module (VRM): Before any planning occurs, the VRM analyzes an image of the scene paired with the PDDL domain specification to foresee downward refinement failures and infer general spatial constraints to prevent them. These are specified as constraints on the SMT solver.
  • Stream-based Grounding: We ground our task plans with the stream-based planning used PDDLStream [4] and COAST [9].
We first briefly introduce the task planning module, then we describe the visual reasoning module in detail, and how it interacts with the downstream task and motion planning components. Our primary contribution lies in the VRM’s ability to foresee and prevent downward refinement failures.
IV-1 Task planning
VIZ-COAST employs a general-purpose task planner based on the Z3 Satisfiability Modulo Theories (SMT) solver [8]. We use the Madagascar [7] encoding for SATPlan [6]. This allows us to define declarative SMT constraints on allowable actions and predicates through Z3’s Python API, taking the form of arbitrary expressions of first-order logic. Other popular planners such as Fast Downward [21] do not have a flexible interface for constraint specification, and require editing the PDDL files directly, introducing additional unnecessary complexity, especially due to the translation step to SAS+ [36] required.
IV-2 Visual reasoning with VLMs
VIZ-COAST’s visual reasoning module (VRM) leverages the common-sense visual and spatial reasoning capabilities of VLMs to foresee potential downward refinement failures and infer necessary planning constraints in advance. As shown in Fig.˜2, this module takes four inputs, which condition each step of prompting: symbolic specifications of the problem and domain in PDDL, an image of the initial state of the scene, a text-based representation of the scene, and an example of a compliant output. Its output is a Python program which encodes the necessary task-level planning constraints to avoid downward refinement failures.
The VRM consists of a 4-step prompting procedure, as described in Fig.˜3. The first step is scene interpretation: from the inputs, generate a description of the scene in natural language. Intuitively, this forces the VLM to interpret the scene in a semantically meaningful way. In the second step, the VLM is asked to identify potential downward refinement issues and describe the necessary constraints to mitigate them in natural language. Then, the third prompt asks the VLM to encode these constraints formally as a Python program, using Z3’s API. The final step is proofreading, in which we simply ask the model to examine the output and ensure it is syntactically compliant and semantically consistent with the natural language description produced by the second prompt. A paraphrased example of a constraint script generated by the VRM which blocks items from being picked out of or placed into closed containers is shown in Fig.˜4.
IV-3 Motion Planning
VIZ-COAST integrates with the COAST framework to perform downward refinement, which itself builds upon PDDLStream [4]’s use of streams for sampling continuous action parameters. As the purpose of our approach is to impose task-level constraints to prevent motion planning failures, we only make modifications at the task planning level. For motion planning, we adopt COAST’s stream-based approach directly without modification.
VIZ-COAST’s goal is to infer the necessary constraints zero-shot, before any planning failures occur. However, in the event that downward refinement still fails, we simply re-run the prompting procedure.
V Experiments
V-A Domains
We evaluated VIZ-COAST on two TAMP domains, each involving a robot arm manipulating objects in a simulated PyBullet environment [37]:
  • Blocks: Rearrange a set of blocks from an initial configuration within a 3x3 grid such that the red block is in the center.
  • Containers: Place items inside closed containers to achieve some goal configuration.
Figure 5: The Block domain (left) requires the robot to rearrange blocks to move the red block to the center tile. The end-effector may only approach the grid from a single direction. The Containers domain (right) requires the robot to place the items in target containers, removing and replacing their lids as needed.
Each domain has a corresponding PDDL description for task planning. Importantly, this symbolic description neglects key considerations for downward refinement, which manifest as motion grounding failures. For each domain, we describe the failure modes and demonstrate that VIZ-COAST infers constraints that not only mitigate these failures more efficiently than prior approaches, but also generalize to the entire class of problem instances posed by the domain.
V-B Baselines
VIZ-COAST enables faster planning than prior state-of-the-art through reduced replanning, as well as higher success rates. We demonstrate this by evaluating VIZ-COAST against 2 strong baselines as well as a naive VLM-based approach:
  • COAST+FD: The original COAST system, which used the Fast Downward task planner.
  • COAST+Z3: A novel baseline we developed, adapting COAST to interface with the SMT-based planner used in VIZ-COAST.
  • VLM-only: A naive baseline which prompts the VLM to directly generate the task plan, given the PDDL domain and problem files and an image of the initial state. Upon failure, VLM-only reprompts.
Both COAST+FD and COAST+Z3 employ COAST’s approach of adding constraints upon encountering downward refinement failures. We use GPT-5 [38] as the VLM in VIZ-COAST’s VRM in experiments and as the VLM-only baseline, although any VLM can be used.
Figure 6: Planning times, refinement failures, and success rates for the Blocks domain, averaged over up to 5 trials. VIZ-COAST was prompted once on the first 1-block problem instance, producing constraints which generalized to all 35 unique problem instances containing up to 7 blocks. By preventing refinement failures and thereby avoiding replanning, VIZ-COAST achieves lower planning times than COAST+FD and COAST+Z3. The VLM-only baseline is far slower than all other approaches, and fails completely on large problem instances.
Figure 7: Planning times, refinement failures, and success rates for the Containers domain, averaged over up to 5 trials. VIZ-COAST was prompted once on the first 1-goal problem instance, and then reprompted on subsequent failures. VIZ-COAST is able to produce successful plans for all problem instances, whereas COAST+FD fails on all instances and COAST+Z3 only succeeds for some 1-goal instances. These failures arise because the task-level specification lacks geometric intuition—the precondition of removing the lid is not captured in the symbolic PDDL description.
V-C Block
The Block domain features a $3\times 3$ grid of tiles on which blocks are placed upright. In each problem instance, number of blocks and their initial configuration may vary, but the goal is always to place the single red block on the middle tile. When picking and placing blocks, the robot may only approach from a single direction; its end effector must be facing away from its base and towards the target block, as shown in Fig.˜5.
Because the robot’s approach vector is restricted, any block obstructing the target along the direction of approach will result in a collision. Therefore, constraints must be generated to disallow pick and place actions if an approach corridor is not clear; this constraint is not encoded within the domain and can only be recovered by reasoning or refinement failure.
Results for the Blocks domain are shown in Fig.˜6. Each method was evaluated on problem instances of up to 7 blocks, with 5 trials per number of blocks, for a total of 35 unique instances. Evaluation proceeded sequentially from 1-block to 7-block instances. Planning time is how long the task planner took to produce a successful high-level plan, and does not include motion planning time, which varied little between methods. Ref. Failures refers to the total number of times each method replanned due to a downward refinement failure. If for a particular problem instance, a method exceeded 40 refinement failures, the instance was deemed a failure. All numbers are averaged over 5 trials. To ensure fair comparison, only problem instances for which at least 3 out of 4 methods succeeded are included in the average. VIZ-COAST’s VRM constraint generation pipeline is called a single time on an example scene extracted from the first single-block problem instance. Constraints persist over trials until a motion grounding failure occurs, at which point the constraint generation pipeline is re-run.
Our experiments show that on the Blocks domain, VIZ-COAST produces successful constraints on the initial call, before any planning occurs, which generalize to all subsequent trials over all block counts. In other words, VIZ-COAST does not encounter any downward refinement failures on the Blocks domain. It is also worth noting that the problem instance from which the constraints were inferred only had one block, yet VIZ-COAST was still able to produce constraints that generalized to several blocks.
In contrast, the COAST-based baselines often encountered 5 or more refinement failures before finding a successful plan on instances with several blocks, resulting in longer total planning times. Interestingly, the VLM-only baseline performed very poorly, succeeding at less than half of all instances, with a 0% success rate on instances with 5 or more blocks.
V-D Containers
The Containers domain consists of 3 containers, each with a movable lid, 6 movable items, and a plate. The robot is able to move the items and lids. In contrast with Blocks, Containers has a fixed initial state, instead varying the goal state between problem instances. Lids obstruct placing items into containers. This domain’s core challenge is that the PDDL domain specification lacks key geometric intuition: removing the lid is not explicitly expressed as a precondition for placing objects into containers. Constraints are required to ensure that objects can only be placed inside or picked out when the target container is open.
Results for the Containers domain are shown in Fig.˜7. Each method was evaluated on problem instances of up to 6 goal predicates with 5 trials per number of goals for a total of 30 unique instances. Goal predicates are randomly generated, each assigning a specific item or lid to be in or on a specific container. VIZ-COAST’s constraint generation pipeline is initially called once before any planning occurs on an example scene extracted from the first problem instance. Upon encountering a downward refinement failure, a new example scene is extracted from the failed instance, and the VRM is called again. As in Blocks, the planning time is how long the task planner took to produce a high-level plan. Ref. Failures refers to the total number of times the method replanned due to a downward refinement failure. For this domain, we report the sum total number of refinement failures for each goal count. If for a particular problem instance a method failed to find a plan within one hour it was deemed a failure. All numbers are averaged over up to 5 trials. Note that COAST+FD failed to find a plan for any problem instance, and COAST+Z3 only succeeded on problem instances with one goal predicate. For goal counts with a 0% success rate, we cannot report planning times or refinement failures.
V-E Robustness
Method Success(%) Total ref. fails 5-block plan time(s)
VIZ-COAST 100 $\textbf{0.5}\pm\textbf{0.878}$ $0.328\pm 0.131$
Condensed 100 $2.0\pm 3.04$ $\textbf{0.296}\pm\textbf{0.090}$
COAST-FD $92.5$ $147.5\pm 104$ $1.257\pm 0.377$
COAST-Z3 8585 $126.5\pm 77.1$ $1.906\pm 0.404$
Table I: Comparison of success rates, average planning times, and average total downward refinement failure counts over 6 sequential 20-instance trials on the Block domain. VIZ-COAST’s 4-stage prompting procedure reliably generalizes across instances, averaging 4 times fewer refinement failures than Condensed (the single-prompt ablation of VIZ-COAST). In contrast, COAST-based baselines encounter frequent refinement failures, resulting in longer planning times.
We conduct an additional comparison in which we run 6 independent trials over 20 sequential problem instances in the Blocks domain for each method. To justify the VRM’s 4-stage prompting procedure, we introduce Condensed, an ablation of VIZ-COAST which produces constraints in just a single prompt from the same inputs.
Of the 6 trials, 3 consist of only 5-block instances, while the other 3 consist of 1 to 4-block instances. VIZ-COAST’s constraint generation pipeline is initially called once before any planning occurs on an example scene extracted from the first problem instance. Upon encountering a downward refinement failure, a new example scene is extracted from the failed instance, and the VRM is called again. Constraints persist between subsequent problem instances. We record each method’s success rate, average planning time on 5-block instances, and the average total number of failures encountered across all 20 instances in a trial. We note that this is not the number of refinement failures encountered on a single problem instance.
As shown in Tbl.˜I, VIZ-COAST’s constraints successfully generalize to all 20 problem instances on the first attempt, prior to any planning, on 4 out of 6 trials. On average, VIZ-COAST encounters $0.5$ motion grounding failures during each 20-instance trial before inferring generalizable constraints. In all trials for both VIZ-COAST and Condensed, any failures encountered were on the first problem instance.
Both VIZ-COAST and Condensed successfully produce downward-refinable task plans for all instances, taking $\sim$0.3 seconds for 5-block instances on average. However, Condensed encounters two motion grounding failures on average per trial, 4 times that of VIZ-COAST, indicating that our 4-stage prompting procedure enables more robust generalization.
Both COAST-based baselines are unable to solve some problem instances, and encounter significantly more downward refinement failures on average across trials. As a result, these methods take 3 to 4 times longer to solve 5-block instances on average, with VIZ-COAST providing a 74% speed-up over our fastest non-ablation baseline, COAST-FD.
VI Discussion
VI-A The Role of VLMs in TAMP
A key distinction between the Blocks and Containers domains is evidenced by the performance characteristics of the VLM-only baseline. In the Containers domain, the baseline successfully solved most problem instances across varying goal counts within a few iterations. This contrasts starkly with its performance in the Blocks domain, where it solved fewer than half of the problem instances, with success rates decreasing drastically as the number of blocks increased. We attribute this discrepancy to fundamental differences in the structure of the task abstraction of the two domains. Even with correct constraints, the Blocks domain requires long-horizon, non-ergodic search to find feasible plans, whereas in the Containers domain, once the key implications on action validity imposed by the container lids is captured (that is, objects cannot be placed in or picked from closed containers), successful plans tend to follow relatively simple patterns of pick-and-place. These results highlight both the strengths and limitations of VLMs in TAMP applications. The VLM-only baseline’s success in the Containers domain demonstrates that VLMs excel at leveraging common-sense visual reasoning to anticipate and avoid downward-refinement failures; conversely, their poor performance in the Blocks domain reveals their limitations in tasks that resolve search. By using VLMs exclusively for inference and imposition of constraints while delegating planning to a TAMP solver, VIZ-COAST exploits VLMs’ strengths while circumventing their inherent weaknesses.
VI-B The Advantage of Constraints in Task Planning
A primary factor contributing to the high failure rates of COAST-based baselines in the Containers domain is the requirement to introduce additional entities and predicates into the PDDL description to express constraints. This modification incurs significant overhead in search (due to the size of the grounding), as well as in the translation step to convert PDDL to SAS+ used by Fast Downward. In contrast, VIZ-COAST encodes constraints directly through the Z3 API, enabling the planner to utilize the original PDDL description while introducing minimal additional complexity. In the Containers domain, this overhead frequently determined whether a method could find a successful plan within seconds or fail entirely. This observation illustrates the computational advantage of SMT-based constraints while simultaneously highlighting a limitation of failure-driven constraint discovery: the process of encountering failures may itself be intractable.
VII Conclusion
We have presented VIZ-COAST, a TAMP solver that uses VLMs’ common-sense geometric reasoning capabilities to constrain TAMP problems a priori. Unlike prior works, the constraints produced by the VLM generalize to unseen problem instances within a broader domain and incur minimal additional overhead by directly interfacing through APIs to an SMT-based task planner. These constraints eliminate potential downward refinement failures that arise between the symbolic abstraction and geometry of the world—preventing the planner from attempting to pick obstructed objects, place objects in cluttered locations, and perform other infeasible actions. We have demonstrated that VIZ-COAST effectively leverages VLMs’ insights to improve planning times by reducing or even eliminating replanning due to downward refinement failures. For example, in the Blocks domain, VIZ-COAST produced general constraints on the first attempt, resulting in task plans that were refinable on the first attempt for all 35 problem instances. In the Containers domain, VIZ-COAST produces constraints that enable solutions to all 30 instances, while baseline methods fail to find solutions due to the complexity of the unconstrained search space.
Overall, VIZ-COAST demonstrates that VLMs can provide valuable insights into problem structure that accelerate search performance. These results are promising for future applications that extract even richer structural information, operate in online settings, or require constraint insertion based on natural language instructions grounded to the physical world. One limitation of our approach is that VLM inference introduces unpredictable latencies. Our method is agnostic to the specific VLM model used; as models become faster and more efficient, we expect this variance to decrease. Evaluating our methodology with local or distilled models represents an interesting direction for future work. We also evaluated our approach exclusively in simulation environments. While we believe that the generality of large pre-trained VLMs will enable constraint extraction in real-world settings, empirical validation of this capability remains future work.

### v2

@@
 We propose VIZ-COAST, a method of leveraging the common-sense spatial reasoning of large pretrained Vision-Language Models to identify issues with downward refinement a priori, bypassing the need to fix these failures during planning.
-Experiments on two challenging TAMP domains show that our approach is able to extract plausible constraints from images and domain descriptions, drastically reducing planning times and, in some cases, eliminating downward refinement failures altogether, generalizing to a diverse range of instances from the broader domain.
+Experiments on three challenging TAMP domains show that our approach is able to extract plausible constraints from images and domain descriptions, drastically reducing planning times and, in some cases, eliminating downward refinement failures altogether, generalizing to a diverse range of instances from the broader domain.
 I Introduction
-Task and Motion Planning (TAMP) [1] offers a framework for solving long-horizon robotic problems that couples symbolic decision-making with geometric feasibility, such as clearing a cluttered table or navigating between rooms where a key must first be found to open a door.
-These problems are compelling not only because completing tasks requires multiple steps, but also because they demand reasoning across different levels of abstraction.
+Task and Motion Planning (TAMP) [garrett2021integrated] offers a framework for solving long-horizon robotic problems that couples symbolic decision-making with geometric feasibility, such as clearing a cluttered table or navigating between rooms where a key must first be found to open a door.
+These problems are compelling not only because tasks require multiple steps, but also because they demand reasoning across different levels of abstraction.
 The robot must determine what actions to take and how to execute them.
@@
 This abstraction allows the planner to efficiently reason about possible sequences of actions without searching in continuous geometric space.
-However, the core difficulty is that these abstractions often cannot express critical geometric information, violating the downward refinement property [2, 3]: a valid task plan may not necessarily correspond to a sequence of feasible motion plans.
+However, the core difficulty is that these abstractions often cannot express critical geometric information, violating the downward refinement property [bacchus1991downward, bacchus1994downward]: a valid task plan may not necessarily correspond to a sequence of feasible motion plans.
 For instance, a symbolic plan may prescribe grasping an object, but that object may be obstructed by another object, preventing any collision-free grasp.
 Figure 1: VIZ-COAST uses VLMs and SMT-based task planning to constrain TAMP problems.
-Our constraints generalize to unseen problem instances within a broader domain zero-shot.
+Our constraints generalize to unseen problem instances within a broader domain.
 VIZ-COAST introduces minimal search overhead, enabling more efficient planning than prior state-of-the-art.
-PDDLStream [4] is a TAMP framework which circumvents downward refinement failures by employing a sample-first approach, in which feasible continuous action parameterizations, such as grasp poses and trajectories, are sampled prior to planning via streams, allowing task planning to be done with geometric feasibility pre-certified.
-However, PDDLStream expends significant effort sampling parameterizations of action groundings which may not be used in the final plan.
-This can become highly inefficient, as the number of possible action groundings grows exponentially with the number of objects.
+PDDLStream [garrett_pddlstream_2020] is a TAMP framework which circumvents downward refinement failures by employing a sample-first approach, in which feasible continuous action parameterizations, such as grasp poses and trajectories, are sampled prior to planning via streams, allowing task planning to be done with geometric feasibility pre-certified.
+However, PDDLStream expends significant effort sampling action groundings which may not be used in the final plan.
+This can be very inefficient, as the number of possible action groundings grows exponentially with the number of objects.
 Alternatively, many TAMP methods take a plan-first approach, creating a complete task plan before attempting downward refinement.
@@
 For example, upon failing to find a feasible grasp pose due to an obstructing object, a constraint may be added to enforce that the offending object must not be in that position when attempting a grasp.
-IDTMP [5] implements this by formulating planning as with SATPlan [6, 7] through Z3 [8], a Satisfiability Modulo Theories (SMT) solver, incrementally adding constraint clauses upon failure.
-COAST [9], a plan-first successor to PDDLStream, adds auxiliary predicates and entities to directly encode constraints in the PDDL description.
+IDTMP [dantam2018incremental] implements this by formulating planning with SATPlan [kautz2006satplan, rintanen2014madagascar] through Z3 [de2008z3], a Satisfiability Modulo Theories (SMT) solver, incrementally adding constraint clauses upon failure.
+COAST [vu2024coast], a plan-first successor to PDDLStream, adds auxiliary predicates and entities to directly encode constraints in the PDDL description.
 All these existing constraint-based approaches ultimately require expending additional search effort to encounter downward-refinement failures and discover the necessary constraints, often replanning several times before finding a plan that can be successfully ground to motion.
 Our goal is to avoid expending this effort by adding the necessary constraints before planning even begins.
-Recently, large pretrained vision language models (VLMs) have demonstrated common-sense geometric and physical reasoning capabilities [10, 11], allowing them to generalize zero-shot to a wide range of settings.
-For this reason, they have been employed for various facets of TAMP, including subgoal generation [12], encoding open-world preferences [13], and correcting errors [14].
-In this work, we propose Visual Insight for Z3 [8]-based Constraints and Streams (VIZ-COAST), a novel approach to TAMP which leverages VLMs to foresee potential downward-refinement failures and add constraints to avoid them before planning begins, eliminating the need to encounter these failures during planning and expend effort replanning.
+Recently, large pretrained vision language models (VLMs) have demonstrated common-sense geometric and physical reasoning capabilities [gao2024physically, chen2024spatialvlm], allowing them to generalize zero-shot to a wide range of settings.
+For this reason, they have been employed for various facets of TAMP, including subgoal generation [yang_guiding_2025], encoding open-world preferences [kumar_open-world_2025], and correcting errors [duan2024aha].
+In this work, we propose Visual Insight for Z3 [de2008z3]-based Constraints and Streams (VIZ-COAST), a novel approach to TAMP which leverages VLMs to foresee potential downward-refinement failures and add constraints to avoid them before planning begins, eliminating the need to encounter these failures during planning and expend effort replanning.
+In challenging domains where initial constraints do not account for all failure modes, VIZ-COAST analyzes multiple problems to revise constraints using feedback from downstream planning in a closed-loop manner.
 As shown in Fig.˜1, we design a pipeline that guides a VLM to produce constraints that directly interface with an SMT-based task planner.
 In contrast to COAST, whose auxiliary PDDL constructions introduce significant search overhead, the SMT-based solver allows us to seamlessly incorporate constraints while using the original PDDL description.
-Through experiments on two challenging TAMP domains, we demonstrate that VIZ-COAST significantly decreases planning time by reducing or even eliminating replanning caused by downward refinement failures.
-We also show that our constraints consistently generalize zero-shot to unseen problem instances within a broader domain.
+Through experiments on three challenging TAMP domains, we demonstrate that VIZ-COAST significantly decreases planning time by reducing or even eliminating replanning caused by downward refinement failures.
+We also show that our constraints consistently generalize to unseen problem instances within a broader domain.
 II Background and Related Work
-Figure 2: The VIZ-COAST architecture.
-Our Visual Reasoning Module takes as input an example scene, consisting of an image and the geometric state, the PDDL domain description, and an example of a syntactically compliant output.
-It produces a Python file encoding constraints through a high-level API to the SMT-based task planner.
-When presented with a new problem instance, the task planner takes these constraints as input, in addition to the PDDL domain and new problem descriptions.
-The planner applies the constraints to block geometrically infeasible task plans.
-The high-level plan produced is then grounded by a streams-based motion planner to produce an executable continuous motion plan.
-In this section, we begin with a general overview of modern task and motion planning research and its main challenges in Sec.˜II-A.
-We then review prior approaches in Sec.˜II-B that accelerate TAMP by constraining the search space, highlighting the need for a generalizable method capable of pruning before planning begins.
+We begin with a general overview of modern task and motion planning and its main challenges in Sec.˜II-A.
+We then review prior approaches in Sec.˜II-B that accelerate TAMP by constraining the search space, highlighting the need for a generalizable method capable of pruning before planning.
 Lastly, in Sec.˜II-C, we explain how VLMs’ spatial reasoning capabilities make such a method possible and clarify how our approach differs from prior work.
 II-A Task and Motion Planning
-Task and motion planning (TAMP) decomposes continuous long-horizon motion planning into search in both discrete (task-level) and continuous (motion-level) spaces [1].
-The task level planning is done over an abstraction of the robot’s capabilities, and produces a sequence of symbolic actions to transition from a given start state to a desired goal condition (e.g., STRIPS [15]).
-In this work we formalize task planning using Planning Domain Definition Language (PDDL) [16] due to its expressive power and widespread adoption.
-A core challenge in TAMP is downward-refinement failure [2, 3], which occurs when a symbolic task plan becomes infeasible at the motion level.
-PDDLStream [4] addresses this issue by taking a sample-first approach, performing continuous sampling before task planning via declarative streams, allowing earlier reasoning about geometric feasibility.
-Other works [17, 18, 19, 9] take a plan-first approach, designing algorithms to add constraints upon encountering failures which inform task-level replanning to prevent revisiting infeasible options.
+Task and motion planning (TAMP) decomposes continuous long-horizon motion planning into search in both discrete (task-level) and continuous (motion-level) spaces [garrett2021integrated].
+The task level planning is done over an abstraction of the robot’s capabilities, and produces a sequence of symbolic actions to transition from a given start state to a desired goal condition (e.g., STRIPS [fikes1971strips]).
+In this work, we formalize task planning using Planning Domain Definition Language (PDDL) [mcdermott1998pddl, fox2003pddl2] due to its expressive power and widespread adoption.
+A core challenge in TAMP is downward-refinement failure [bacchus1991downward, bacchus1994downward], which occurs when a symbolic task plan becomes infeasible at the motion level.
+PDDLStream [garrett_pddlstream_2020] addresses this issue by taking a sample-first approach, performing continuous sampling before task planning via declarative streams, allowing earlier reasoning about geometric feasibility.
+Other works [kaelbling_hierarchical_2011, srivastava2014combined, dantam2013motion, vu2024coast] take a plan-first approach, designing algorithms to add constraints upon encountering failures which inform task-level replanning to prevent revisiting infeasible options.
 In the next section, we expand on the role of constraints and other methods of reducing the search space in TAMP.
@@
 The purpose of constraints in both IDTMP and COAST is to prune the task-level search space to avoid repeating failures, though they differ in implementation.
-COAST [9] is a plan-first adaptation of PDDLStream which encodes constraints by modifying the PDDL description, adding auxiliary predicates and entities to block actions under specific conditions.
+COAST [vu2024coast] is a plan-first adaptation of PDDLStream which encodes constraints by modifying the PDDL description, adding auxiliary predicates and entities to block actions under specific conditions.
 While this approach allows COAST to be planner-agnostic, encoding constraints directly in PDDL limits the constraints that may be expressed, and the auxiliary constructions introduce significant search overhead.
-PDDL 3.0 [20] introduced constraints on valid states, but is poorly supported by existing planners (e.g., Fast Downward [21]) and cannot be reliably generated by LLMs due to limited training data.
-In contrast, we take the approach done by IDTMP [5], and solve for valid task plans through SATPlan [6, 7] encoded in a Satisfiability Modulo Theories (SMT) solver (in particular, Z3 [8]).
+PDDL 3.0 [Gerevini2005PDDLConstraints] introduced constraints on valid states, but is poorly supported by existing planners (e.g., Fast Downward [helmert2006fast]) and cannot be reliably generated by LLMs due to limited training data.
+In contrast, we take the approach used by IDTMP [dantam2018incremental], and solve for valid task plans through SATPlan [kautz2006satplan, rintanen2014madagascar] encoded in a Satisfiability Modulo Theories (SMT) solver (in particular, Z3 [de2008z3]).
 This formulation enables constraints to be directly encoded in the SMT solver through simple Python programming, enabling efficient and flexible constraint specification.
-While heuristic search approaches typically exhibit faster planning times [22], SATPlan encodings offer far greater specification flexibility and can avoid the costly translation overhead in FastDownward [21] for many problems.
-Although existing methods [9, 5] are able to overcome downward refinement issues, they discover constraints only by encountering motion grounding failures during execution.
+While heuristic search approaches typically exhibit faster planning times [taitler20242023], SATPlan encodings offer far greater specification flexibility and can avoid the costly translation overhead in FastDownward [helmert2006fast] for many problems.
+Although existing methods [vu2024coast, dantam2018incremental] are able to overcome downward refinement issues, they discover constraints only by encountering motion grounding failures during execution.
 This reactive process is often repeated for several iterations before uncovering the necessary constraints to solve a particular problem instance, and cannot generalize to unseen instances.
 Alternative approaches have explored predicting action feasibility to avoid refinement failures.
-Wells et al.
-[23] learn feasibility constraints for tabletop manipulation through domain-specific training, while Xu et al.
-[24] employ neural feasibility checks to identify infeasible actions before motion planning.
-Driess et al.
-[25] use deep visual heuristics to assess feasibility in mixed-integer manipulation planning, and [26] leverage visual reasoning to predict valid actions for TAMP.
+wells2019learning learn feasibility constraints for tabletop manipulation through domain-specific training, while xu2022accelerating employ neural feasibility checks to identify infeasible actions before motion planning.
+driess2020deep use deep visual heuristics to assess feasibility in mixed-integer manipulation planning, and [20-driess-RSS] leverage visual reasoning to predict valid actions for TAMP.
 Although promising, these methods require extensive domain-specific training data and may not generalize well to new environments or object configurations.
@@
 II-C Vision Language Models for TAMP
-LLMs have demonstrated that they are suitable for high [27].
-VLMs further integrate visual and textual inputs to enable scene understanding [28], spatial reasoning [11], and grounded action execution [29, 14].
+LLMs have demonstrated that they are suitable for high-level reasoning [aghzal2025survey].
+VLMs further integrate visual and textual inputs to enable scene understanding [shek2024lancar], spatial reasoning [chen2024spatialvlm], and grounded action execution [liu2024vision, duan2024aha].
 These capabilities make VLMs especially promising for addressing the geometric reasoning challenges in TAMP.
 Several works have integrated LLMs and VLMs into TAMP pipelines for various purposes.
-Zhang et al.
-[30] use VLMs to ground task planners by validating symbolic predicates and providing incremental feedback.
-Hu et al.
-[31] uses a VLM to generate physically feasible actions.
-ReplanVLM [32] employs VLMs for plan generation with replanning upon grounding failures.
-CaStL [33] translates natural language task descriptions into PDDL with constraints reflecting user preferences.
-OWL-TAMP [13] encodes open-world preferences as constraints on symbolic and geometric states.
+zhang2023grounding use VLMs to ground task planners by validating symbolic predicates and providing incremental feedback.
+hu2023look uses a VLM to generate physically feasible actions.
+ReplanVLM [mei2024replanvlm] employs VLMs for plan generation with replanning upon grounding failures.
+CaStL [guo_castl_2025] translates natural language task descriptions into PDDL with constraints reflecting user preferences.
+OWL-TAMP [kumar_open-world_2025] encodes open-world preferences as constraints on symbolic and geometric states.
 Both CaStL and OWL-TAMP operate at the problem specification level, aiming to capture preferences rather than prevent downward refinement failures.
-VLM-TAMP [12] uses VLMs’ common-sense knowledge to generate intermediate subgoals for long-horizon planning; however, it assumes plans can always be refined at the motion level, making it prone to failure when this assumption fails.
-CLIMB [34] incrementally updates the domain description upon failure with an LLM to continually improve task performance; however, it is restricted to constraints expressible in PDDL, and does not make use of visual input.
+VLM-TAMP [yang_guiding_2025] uses VLMs to generate intermediate subgoals for long-horizon planning.
+While this accelerates task-level plan search, it relies on the assumption of downward refinability.
+CLIMB [byrnes_climb_2025] incrementally updates the domain description upon failure with an LLM to continually improve task performance; however, it is restricted to constraints expressible in PDDL, and does not make use of visual input.
 Our work leverages VLMs’ spatial reasoning capabilities to proactively identify general geometric constraints that prevent downward refinement failures before planning begins.
@@
 III Problem Formulation
-Our work investigates the use of VLMs to accelerate solving the TAMP problem.
-In particular, our method is capable of producing constraints which generalize to a range of problem instances within a broader domain.
-When inferring constraints for a particular domain, we assume access to a model of the robot and its environment, as well as a Planning Domain Definition Language (PDDL) [35] domain description which specifies the available actions, predicates, and possible objects.
-Additionally, we assume access to an image of some example scene within the domain.
-We assume knowledge of the underlying geometric state of this example scene, consisting of the coordinates and type of each object present.
-The constraint inference step does not involve a PDDL problem description.
-Once constraints have been generated, they are used to plan on specific problem instances.
-At this stage, the specific PDDL problem description is accessed by the task planner.
+Our work investigates the use of VLMs to accelerate TAMP. In particular, our method is capable of producing constraints which generalize to a range of problem instances within a broader domain.
+When inferring constraints for a particular domain, we assume access to a model of the robot and its environment, as well as a symbolic Planning Domain Definition Language (PDDL) [mcdermott1998pddl] domain description.
+Additionally, we assume access to some example scene within the domain, consisting of several components: (1) a symbolic PDDL problem description; (2) the geometric coordinates and dimensions of each object present; (3) an image of the initial state of the scene which clearly shows all objects.
+Lastly, we assume the set of possible unique object instances is finite and known.
+Once constraints have been generated, they are used to plan on subsequent problem instances.
 We use the PDDL standard for symbolic task planning.
-For a given domain, a PDDL domain specifies the allowed types, predicates and actions.
+For a given domain, the PDDL domain description specifies the allowed types, predicates and actions.
 Actions are defined by their preconditions, expressed as logical formulas over predicates, and their effects, which specify how the truth values of a subset of predicates are updated in the next state.
 An instance of a problem in a particular domain defines a set of objects, their initial state, and the goal conditions.
-States are represented as a set of predicates which true.
-We adopt the closed-world assumption, in predicates not specified to be true are implicitly assumed to be false.
+States are represented as the set of predicates which are true.
+We adopt the closed-world assumption, in which predicates not specified to be true are implicitly assumed to be false.
 The problem of task planning is to find a sequence of actions which achieves the goal from the initial state.
@@
 IV Methodology
-Figure 3: VIZ-COAST’s Visual Reasoning Module infers constraints through a 4-step prompting procedure.
-First, it is asked to provide an interpretation of the scene based on an image of the example problem instance and the PDDL domain description.
-Then, it is asked to articulate in natural language the necessary constraints.
-After that, the VLM is asked to formally encode the constraints it identified in Python using Z3’s API, referencing a structural example.
-Finally, it proofreads the script to ensure syntactic compliance.
+Figure 2: The VIZ-COAST architecture.
+Our Visual Reasoning Module takes as input an example scene, consisting of an image and the geometric state, the PDDL domain and problem description, and the grasp information in natural language.
+It produces a Python file encoding constraints through a high-level API to the SMT-based task planner.
+When presented with a new problem instance, the task planner takes these constraints as input, in addition to the PDDL domain and new problem descriptions.
+The planner applies the constraints to block geometrically infeasible task plans.
+The high-level plan produced is then grounded by a streams-based motion planner to produce an executable continuous motion plan.
 Figure 3: Paraphrased example of a constraints file produced by the Visual Reasoning Module for the Containers domain.
@@
 VIZ-COAST consists of 3 primary components: The task planning module, the visual reasoning module (VRM), and the motion grounding module.
-  • SMT-based Task planner: We use an Satisfiability Module Theories (SMT) solver—Z3 [8], in particular—that uses a SATPlan encoding [6, 7] to enable flexible constraint specification while introducing minimal search overhead.
-This is similar to the task planning framework used in IDTMP [5] and CaStL [33].
-  • Visual Reasoning Module (VRM): Before any planning occurs, the VRM analyzes an image of the scene paired with the PDDL domain specification to foresee downward refinement failures and infer general spatial constraints to prevent them.
-These are specified as constraints on the SMT solver.
-  • Stream-based Grounding: We ground our task plans with the stream-based planning used PDDLStream [4] and COAST [9].
-We first briefly introduce the task planning module, then we describe the visual reasoning module in detail, and how it interacts with the downstream task and motion planning components.
+  • SMT-based Task planner: We use a Satisfiability Modulo Theories (SMT) solver—Z3 [de2008z3], in particular—that uses a SATPlan encoding [kautz2006satplan, rintanen2014madagascar] to enable flexible constraint specification while introducing minimal search overhead.
+This is similar to the task planning framework used in IDTMP [dantam2018incremental] and CaStL [guo_castl_2025].
+  • Visual Reasoning Module (VRM): Before planning begins, the VRM uses a VLM to analyze an example scene to foresee downward refinement failures and infer general constraints to prevent them, encoded as constraints on the SMT solver.
+  • Stream-based Grounding: We ground task plans to motion with the streams framework used in PDDLStream [garrett_pddlstream_2020] and COAST [vu2024coast].
+We first briefly introduce the task planning module, then we describe the VRM in detail, and how it interacts with the downstream task and motion planning components.
+The full architecture is shown in Fig.˜2.
 Our primary contribution lies in the VRM’s ability to foresee and prevent downward refinement failures.
 IV-1 Task planning
-VIZ-COAST employs a general-purpose task planner based on the Z3 Satisfiability Modulo Theories (SMT) solver [8].
-We use the Madagascar [7] encoding for SATPlan [6].
+VIZ-COAST employs a general-purpose task planner based on the Z3 SMT solver [de2008z3].
+We use the Madagascar [rintanen2014madagascar] encoding for SATPlan [kautz2006satplan].
 This allows us to define declarative SMT constraints on allowable actions and predicates through Z3’s Python API, taking the form of arbitrary expressions of first-order logic.
-Other popular planners such as Fast Downward [21] do not have a flexible interface for constraint specification, and require editing the PDDL files directly, introducing additional unnecessary complexity, especially due to the translation step to SAS+ [36] required.
+Other popular planners such as Fast Downward [helmert2006fast] do not have a flexible interface for constraint specification, and require editing the PDDL files directly, introducing additional unnecessary complexity, especially due to the translation step to SAS+ [jonsson1998state] required.
 IV-2 Visual reasoning with VLMs
 VIZ-COAST’s visual reasoning module (VRM) leverages the common-sense visual and spatial reasoning capabilities of VLMs to foresee potential downward refinement failures and infer necessary planning constraints in advance.
-As shown in Fig.˜2, this module takes four inputs, which condition each step of prompting: symbolic specifications of the problem and domain in PDDL, an image of the initial state of the scene, a text-based representation of the scene, and an example of a compliant output.
+This module takes five inputs, which condition each step of prompting: symbolic specification of the problem and domain in PDDL, an image of the initial state of the scene, a text-based geometric representation of the scene, a grasp specification, and an example of a compliant output.
+The grasp specification is a short natural-language description of how grasps and approach poses are sampled in the downstream streams-based grounding.
 Its output is a Python program which encodes the necessary task-level planning constraints to avoid downward refinement failures.
-The VRM consists of a 4-step prompting procedure, as described in Fig.˜3.
-The first step is scene interpretation: from the inputs, generate a description of the scene in natural language.
-Intuitively, this forces the VLM to interpret the scene in a semantically meaningful way.
-In the second step, the VLM is asked to identify potential downward refinement issues and describe the necessary constraints to mitigate them in natural language.
-Then, the third prompt asks the VLM to encode these constraints formally as a Python program, using Z3’s API.
-The final step is proofreading, in which we simply ask the model to examine the output and ensure it is syntactically compliant and semantically consistent with the natural language description produced by the second prompt.
-A paraphrased example of a constraint script generated by the VRM which blocks items from being picked out of or placed into closed containers is shown in Fig.˜4.
+The VRM consists of an initial 2-step prompting procedure, followed by a feedback mechanism.
+For initial constraint generation, the VLM is first asked to identify potential downward refinement issues and articulate the necessary constraints to mitigate them in natural language.
+Then, the VLM is asked to encode these constraints as a Python program, using Z3’s API.
+A paraphrased example of a constraint script generated by the VRM is shown in Fig.˜3.
+In the event that planning still fails with these initial constraints, the VRM continues prompting the VLM in a multi-turn conversation to update the constraints using feedback from planning.
+There are three failure cases from planning: if the task planner fails to find a plan, the VLM is asked to make the constraints less restrictive; if the constraints raise errors during application, the VLM is given the error message and asked to debug the constraints; if a task plan is produced but results in downward refinement failure, the task plan and the stream that failed to be successfully sampled are written to a feedback cache, a temporary state-action block is added to the constraints to prevent revisiting the failed action, and planning is reattempted.
+If no plan succeeds after $K$ such attempts, the temporary blocks are removed and the VLM is asked to analyze the $K$ feedback entries and revise the constraints to account for these failures.
+This allows the VLM to reason over several unique failures of the current set of constraints to make a more informed correction.
 IV-3 Motion Planning
-VIZ-COAST integrates with the COAST framework to perform downward refinement, which itself builds upon PDDLStream [4]’s use of streams for sampling continuous action parameters.
+VIZ-COAST integrates with the COAST framework to perform downward refinement, which itself builds upon PDDLStream [garrett_pddlstream_2020]’s use of streams for sampling continuous action parameters.
 As the purpose of our approach is to impose task-level constraints to prevent motion planning failures, we only make modifications at the task planning level.
 For motion planning, we adopt COAST’s stream-based approach directly without modification.
-VIZ-COAST’s goal is to infer the necessary constraints zero-shot, before any planning failures occur.
-However, in the event that downward refinement still fails, we simply re-run the prompting procedure.
 V Experiments
 V-A Domains
-We evaluated VIZ-COAST on two TAMP domains, each involving a robot arm manipulating objects in a simulated PyBullet environment [37]:
-  • Blocks: Rearrange a set of blocks from an initial configuration within a 3x3 grid such that the red block is in the center.
+We evaluated VIZ-COAST on three TAMP domains, each involving a robot arm manipulating objects in a simulated PyBullet environment [pybullet]:
+  • Blocks: Rearrange blocks within a 3x3 grid such that the red block is in the center.
   • Containers: Place items inside closed containers to achieve some goal configuration.
-Figure 5: The Block domain (left) requires the robot to rearrange blocks to move the red block to the center tile.
+  • Packing: Pack several items of varying shape into a container such that the lid may be placed on top.
+Figure 4: The Blocks domain (left) requires the robot to rearrange blocks to move the red block to the center tile.
 The end-effector may only approach the grid from a single direction.
-The Containers domain (right) requires the robot to place the items in target containers, removing and replacing their lids as needed.
+The Containers domain (middle) requires the robot to place the items in target containers, removing and replacing their lids as needed.
+The Packing domain (right) requires the robot to cleverly pack all items into a container and close the lid.
 Each domain has a corresponding PDDL description for task planning.
 Importantly, this symbolic description neglects key considerations for downward refinement, which manifest as motion grounding failures.
-For each domain, we describe the failure modes and demonstrate that VIZ-COAST infers constraints that not only mitigate these failures more efficiently than prior approaches, but also generalize to the entire class of problem instances posed by the domain.
+For each domain, we describe the failure modes and demonstrate that VIZ-COAST infers constraints that not only mitigate these failures more efficiently than prior approaches, but also generalize across the class of problem instances posed by the domain.
 V-B Baselines
 VIZ-COAST enables faster planning than prior state-of-the-art through reduced replanning, as well as higher success rates.
-We demonstrate this by evaluating VIZ-COAST against 2 strong baselines as well as a naive VLM-based approach:
+We demonstrate this by evaluating VIZ-COAST against two strong baselines as well as two naive approaches.
   • COAST+FD: The original COAST system, which used the Fast Downward task planner.
   • COAST+Z3: A novel baseline we developed, adapting COAST to interface with the SMT-based planner used in VIZ-COAST.
+  • Incremental: A naive learning-from-failure method which adds a single state-action block to the SMT-based planner upon refinement failure, and iteratively extends task planning horizon, similar to IDTMP [dantam2018incremental].
   • VLM-only: A naive baseline which prompts the VLM to directly generate the task plan, given the PDDL domain and problem files and an image of the initial state. Upon failure, VLM-only reprompts.
-Both COAST+FD and COAST+Z3 employ COAST’s approach of adding constraints upon encountering downward refinement failures.
-We use GPT-5 [38] as the VLM in VIZ-COAST’s VRM in experiments and as the VLM-only baseline, although any VLM can be used.
-Figure 6: Planning times, refinement failures, and success rates for the Blocks domain, averaged over up to 5 trials.
-VIZ-COAST was prompted once on the first 1-block problem instance, producing constraints which generalized to all 35 unique problem instances containing up to 7 blocks.
-By preventing refinement failures and thereby avoiding replanning, VIZ-COAST achieves lower planning times than COAST+FD and COAST+Z3.
-The VLM-only baseline is far slower than all other approaches, and fails completely on large problem instances.
-Figure 7: Planning times, refinement failures, and success rates for the Containers domain, averaged over up to 5 trials.
-VIZ-COAST was prompted once on the first 1-goal problem instance, and then reprompted on subsequent failures.
-VIZ-COAST is able to produce successful plans for all problem instances, whereas COAST+FD fails on all instances and COAST+Z3 only succeeds for some 1-goal instances.
-These failures arise because the task-level specification lacks geometric intuition—the precondition of removing the lid is not captured in the symbolic PDDL description.
-V-C Block
-The Block domain features a $3\times 3$ grid of tiles on which blocks are placed upright.
-In each problem instance, number of blocks and their initial configuration may vary, but the goal is always to place the single red block on the middle tile.
-When picking and placing blocks, the robot may only approach from a single direction; its end effector must be facing away from its base and towards the target block, as shown in Fig.˜5.
+Both COAST+FD and COAST+Z3 employ COAST’s approach for adding constraints upon encountering downward refinement failures.
+We used Gemini 3 Flash on medium reasoning level as the VLM in VIZ-COAST’s VRM and the VLM-only baseline for baseline comparison experiments.
+A comparison of different VLM providers is presented in Section V-F.
+Figure 5: Success rates, refinement failures, and planning times for the Blocks domain.
+VIZ-COAST was prompted once on the first 2-block problem, producing constraints which generalized to all 30 unique problems containing up to 7 blocks.
+VIZ-COAST achieves fewer refinement failures than COAST-based methods, while exhibiting comparable planning times.
+Figure 6: Success rates, planning times, and refinement failures for the Containers domain.
+VIZ-COAST was prompted once on the first problem, producing constraints which generalized to all 30 unique problems containing up to 6 goal predicates.
+VIZ-COAST is able to produce successful plans for all problems, whereas COAST+FD fails on all problems and COAST+Z3 only succeeds for some 1-goal problems.
+Figure 7: Success rates, refinement failures, and planning times for the Packing domain.
+VIZ-COAST was prompted once on the first 3-item problem to produce initial constraints, and then re-prompted on subsequent failures with feedback to revise constraints.
+VIZ-COAST achieves 100% success and fastest planning times on larger problem instances, whereas COAST-based methods become increasingly slower or fail completely.
+The failures of VIZ-COAST’s no-feedback ablation on 5-item problems shows the necessity of closing the constraint generation-planning loop in the VRM.
+V-C Blocks
+The Blocks domain features a $3\times 3$ grid of tiles on which blocks are placed upright.
+In each problem instance, the number of blocks and their initial configuration may vary, but the goal is always to place the single red block on the middle tile.
+When picking and placing blocks, the robot may only approach and grasp from a single direction; its end effector must be facing away from its base and towards the target block, as shown in Fig.˜4.
+This information is given to the VRM as part of the grasp specification.
 Because the robot’s approach vector is restricted, any block obstructing the target along the direction of approach will result in a collision.
 Therefore, constraints must be generated to disallow pick and place actions if an approach corridor is not clear; this constraint is not encoded within the domain and can only be recovered by reasoning or refinement failure.
-Results for the Blocks domain are shown in Fig.˜6.
-Each method was evaluated on problem instances of up to 7 blocks, with 5 trials per number of blocks, for a total of 35 unique instances.
-Evaluation proceeded sequentially from 1-block to 7-block instances.
+Results for the Blocks domain are shown in Fig.˜5.
+Each method was evaluated on problem instances of up to 2-7 blocks, with 5 problems per number of blocks, for a total of 30 unique instances.
+Evaluation proceeded sequentially from 2-block to 7-block instances.
 Planning time is how long the task planner took to produce a successful high-level plan, and does not include motion planning time, which varied little between methods.
@@
 Failures refers to the total number of times each method replanned due to a downward refinement failure.
-If for a particular problem instance, a method exceeded 40 refinement failures, the instance was deemed a failure.
-All numbers are averaged over 5 trials.
-To ensure fair comparison, only problem instances for which at least 3 out of 4 methods succeeded are included in the average.
+If, for a particular problem, a method exceeded 100 refinement failures or a timeout of 600 seconds, the instance was deemed a failure.
+Metrics are averaged over up to all 5 problems at each number of goals.
+To ensure a fair comparison, only data points for which at least 3 methods succeeded were included in the average.
 VIZ-COAST’s VRM constraint generation pipeline is called a single time on an example scene extracted from the first single-block problem instance.
-Constraints persist over trials until a motion grounding failure occurs, at which point the constraint generation pipeline is re-run.
+Constraints persist over trials until a motion grounding failure occurs, at which point the VRM’s failure feedback mechanism updates the constraints.
 Our experiments show that on the Blocks domain, VIZ-COAST produces successful constraints on the initial call, before any planning occurs, which generalize to all subsequent trials over all block counts.
+VIZ-COAST’s feedback mechanism used re-prompt threshold $K=1$, such that any failure would immediately trigger a constraint update.
 In other words, VIZ-COAST does not encounter any downward refinement failures on the Blocks domain.
-It is also worth noting that the problem instance from which the constraints were inferred only had one block, yet VIZ-COAST was still able to produce constraints that generalized to several blocks.
-In contrast, the COAST-based baselines often encountered 5 or more refinement failures before finding a successful plan on instances with several blocks, resulting in longer total planning times.
-Interestingly, the VLM-only baseline performed very poorly, succeeding at less than half of all instances, with a 0% success rate on instances with 5 or more blocks.
+In contrast, COAST-based baselines often encountered as many as 10 failures before finding a successful plan on many-block instances.
+Though COAST-based methods are faster on simpler problems, their planning times show an exponential upward trend as the number of blocks increases.
 V-D Containers
-The Containers domain consists of 3 containers, each with a movable lid, 6 movable items, and a plate.
-The robot is able to move the items and lids.
+The Containers domain consists of 3 containers with removable lids, 6 movable items, and a static plate where lids may be placed.
 In contrast with Blocks, Containers has a fixed initial state, instead varying the goal state between problem instances.
 Lids obstruct placing items into containers.
+Objects are grasped and approached from above.
 This domain’s core challenge is that the PDDL domain specification lacks key geometric intuition: removing the lid is not explicitly expressed as a precondition for placing objects into containers.
 Constraints are required to ensure that objects can only be placed inside or picked out when the target container is open.
-Results for the Containers domain are shown in Fig.˜7.
-Each method was evaluated on problem instances of up to 6 goal predicates with 5 trials per number of goals for a total of 30 unique instances.
-Goal predicates are randomly generated, each assigning a specific item or lid to be in or on a specific container.
-VIZ-COAST’s constraint generation pipeline is initially called once before any planning occurs on an example scene extracted from the first problem instance.
-Upon encountering a downward refinement failure, a new example scene is extracted from the failed instance, and the VRM is called again.
-As in Blocks, the planning time is how long the task planner took to produce a high-level plan.
-Ref.
-Failures refers to the total number of times the method replanned due to a downward refinement failure.
-For this domain, we report the sum total number of refinement failures for each goal count.
-If for a particular problem instance a method failed to find a plan within one hour it was deemed a failure.
-All numbers are averaged over up to 5 trials.
-Note that COAST+FD failed to find a plan for any problem instance, and COAST+Z3 only succeeded on problem instances with one goal predicate.
-For goal counts with a 0% success rate, we cannot report planning times or refinement failures.
-V-E Robustness
-Method Success(%) Total ref. fails 5-block plan time(s)
-VIZ-COAST 100 $\textbf{0.5}\pm\textbf{0.878}$ $0.328\pm 0.131$
-Condensed 100 $2.0\pm 3.04$ $\textbf{0.296}\pm\textbf{0.090}$
-COAST-FD $92.5$ $147.5\pm 104$ $1.257\pm 0.377$
-COAST-Z3 8585 $126.5\pm 77.1$ $1.906\pm 0.404$
-Table I: Comparison of success rates, average planning times, and average total downward refinement failure counts over 6 sequential 20-instance trials on the Block domain.
-VIZ-COAST’s 4-stage prompting procedure reliably generalizes across instances, averaging 4 times fewer refinement failures than Condensed (the single-prompt ablation of VIZ-COAST).
-In contrast, COAST-based baselines encounter frequent refinement failures, resulting in longer planning times.
-We conduct an additional comparison in which we run 6 independent trials over 20 sequential problem instances in the Blocks domain for each method.
-To justify the VRM’s 4-stage prompting procedure, we introduce Condensed, an ablation of VIZ-COAST which produces constraints in just a single prompt from the same inputs.
-Of the 6 trials, 3 consist of only 5-block instances, while the other 3 consist of 1 to 4-block instances.
-VIZ-COAST’s constraint generation pipeline is initially called once before any planning occurs on an example scene extracted from the first problem instance.
-Upon encountering a downward refinement failure, a new example scene is extracted from the failed instance, and the VRM is called again.
-Constraints persist between subsequent problem instances.
-We record each method’s success rate, average planning time on 5-block instances, and the average total number of failures encountered across all 20 instances in a trial.
-We note that this is not the number of refinement failures encountered on a single problem instance.
-As shown in Tbl.˜I, VIZ-COAST’s constraints successfully generalize to all 20 problem instances on the first attempt, prior to any planning, on 4 out of 6 trials.
-On average, VIZ-COAST encounters $0.5$ motion grounding failures during each 20-instance trial before inferring generalizable constraints.
-In all trials for both VIZ-COAST and Condensed, any failures encountered were on the first problem instance.
-Both VIZ-COAST and Condensed successfully produce downward-refinable task plans for all instances, taking $\sim$0.3 seconds for 5-block instances on average.
-However, Condensed encounters two motion grounding failures on average per trial, 4 times that of VIZ-COAST, indicating that our 4-stage prompting procedure enables more robust generalization.
-Both COAST-based baselines are unable to solve some problem instances, and encounter significantly more downward refinement failures on average across trials.
-As a result, these methods take 3 to 4 times longer to solve 5-block instances on average, with VIZ-COAST providing a 74% speed-up over our fastest non-ablation baseline, COAST-FD.
+Results for the Containers domain are shown in Fig.˜6.
+Each method was evaluated on a fixed set of problem instances of 1-6 goal predicates with 5 problems per number of goals for a total of 30 unique instances.
+Each problem’s goal is a randomly generated set of predicates, each assigning a specific item or lid to be in or on a specific container.
+VIZ-COAST was able to find generalizable constraints on the first attempt.
+Although it encountered a small number of refinement failures on subsequent problems, successful plans were found within a constant number ($K=5$) incremental iterations.
+COAST+FD failed to find a plan for any problem instance, and COAST+Z3 only succeeded on problem instances with one goal predicate.
+For goal counts with a 0% success rate, we cannot report planning times or failure count.
+V-E Packing
+The Packing domain consists of 1 container with a lid and 3 kinds of packable items: slabs, big blocks, and small blocks.
+In each problem, all items start on the table, and the goal is to pack them all into the container and then place the lid on top.
+The symbolic domain description does not account for container’s limited internal volume, thus without constraints a trivial task plan is to pack all objects directly on the container floor.
+However, this is often geometrically infeasible as 1 slab occupies the entire floor.
+Constraints must be added to force the task planner to stack items in a configuration such that they will all fit horizontally without exceeding the vertical limit, so that the lid can be placed on top.
+Results for the Packing domain are shown in Fig.˜7.
+Each method was evaluated on a fixed set of problem instances of 3, 4, 5, and 6 packable items where problem counts were 7, 8, 8, and 7, respectively, for a total of 30 unique instances.
+VIZ-COAST’s feedback mechanism used re-prompt threshold $K=5$.
+Our results show that VIZ-COAST consistently achieves the lowest planning times on problem instances with 5 or 6 items to pack.
+In contrast, COAST+FD is fast on easier problems, but quickly begins to slow down and fail as complexity increases, failing completely on 6-item problems.
+COAST+Z3 fails on all problems.
+Remarkably, the VLM-only baseline achieves 100% success and a comparable number of refinement failures to VIZ-COAST, although it takes much longer to plan.
+Unlike in Blocks and Containers, VIZ-COAST was unable to find fully-generalizable constraints on the initial attempt, requiring continued prompting using feedback from subsequent problems to correct constraints.
+Thus, we compare against an ablation of VIZ-COAST which does not use the feedback mechanism, instead starting a new conversation from scratch upon refinement failure.
+This ablation failed on several 5-item problems, and encountered more refinement failures than VIZ-COAST, demonstrating the benefit of using feedback from planning in the VRM.
+V-F VRM Robustness
+VLM Success Prompting(s) Ref. Failures
+Gemini3-Pro-High 100% $94.09\pm 17.24$ ${\bf 0.00}\pm{\bf 0.00}$
+Gemini3-Pro-Med 100% $54.79\pm 34.60$ $0.05\pm 0.22$
+Gemini3-Flash-Med 100% $61.06\pm 11.24$ $0.15\pm 0.65$
+Gemini3-Flash-Low 100% ${\bf 14.94}\pm{\bf 3.33}$ $0.20\pm 0.40$
+GPT5.2-High 100% $113.68\pm 48.56$ $0.35\pm 0.96$
+GPT5.2-Med 100% $53.91\pm 13.07$ $0.60\pm 1.36$
+GPT5.2-Low 100% $29.25\pm 6.09$ $0.70\pm 0.95$
+Qwen3.5-Plus 100% $184.36\pm 129.12$ $1.05\pm 2.58$
+Qwen3.5-Flash $85\%$ $56.93\pm 49.31$ $4.59\pm 5.26$
+Table I: Comparison of success rates, average time-per-prompt, and average total downward refinement failure counts of various VLM models over 20 independent trials of 4 sequential 4-6 block problems in the Block domain.
+Gemini models performed best, with refinement failures decreasing as reasoning level increased.
+To evaluate the consistency with which the VRM infers constraints that successfully prevent downward refinement failure, we conduct an additional comparison in which we run 20 independent trials over 4 sequential problem instances in the Blocks domain for various VLM configurations.
+We evaluated 5 VLM models at various reasoning levels across 3 providers.
+We report each VLM’s success rate (Success), average time per constraint update (Prompting), and the average total number of refinement failures encountered over all problems in a trial (Ref.
+Failures).
+Metrics are averaged over runs in which all problems succeeded.
+We set the VRM’s re-prompting threshold $K=1$ so that refinement failures would immediately trigger a constraint update.
+The Gemini family of models performed best overall, with Gemini-3-Flash-Low yielding fewer refinement failures on average than GPT-5.2-High, despite taking nearly one-tenth of the time on average.
+Increasing chain-of-thought reasoning corresponded with a reduction in refinement failures, indicating that VLMs’ spatial reasoning capabilities improve with reasoning budget.
+All models achieved a 100% success rate except for Qwen-3.5-Flash.
 VI Discussion
 VI-A The Role of VLMs in TAMP
-A key distinction between the Blocks and Containers domains is evidenced by the performance characteristics of the VLM-only baseline.
-In the Containers domain, the baseline successfully solved most problem instances across varying goal counts within a few iterations.
-This contrasts starkly with its performance in the Blocks domain, where it solved fewer than half of the problem instances, with success rates decreasing drastically as the number of blocks increased.
-We attribute this discrepancy to fundamental differences in the structure of the task abstraction of the two domains.
-Even with correct constraints, the Blocks domain requires long-horizon, non-ergodic search to find feasible plans, whereas in the Containers domain, once the key implications on action validity imposed by the container lids is captured (that is, objects cannot be placed in or picked from closed containers), successful plans tend to follow relatively simple patterns of pick-and-place.
-These results highlight both the strengths and limitations of VLMs in TAMP applications.
-The VLM-only baseline’s success in the Containers domain demonstrates that VLMs excel at leveraging common-sense visual reasoning to anticipate and avoid downward-refinement failures; conversely, their poor performance in the Blocks domain reveals their limitations in tasks that resolve search.
-By using VLMs exclusively for inference and imposition of constraints while delegating planning to a TAMP solver, VIZ-COAST exploits VLMs’ strengths while circumventing their inherent weaknesses.
-VI-B The Advantage of Constraints in Task Planning
-A primary factor contributing to the high failure rates of COAST-based baselines in the Containers domain is the requirement to introduce additional entities and predicates into the PDDL description to express constraints.
-This modification incurs significant overhead in search (due to the size of the grounding), as well as in the translation step to convert PDDL to SAS+ used by Fast Downward.
-In contrast, VIZ-COAST encodes constraints directly through the Z3 API, enabling the planner to utilize the original PDDL description while introducing minimal additional complexity.
-In the Containers domain, this overhead frequently determined whether a method could find a successful plan within seconds or fail entirely.
-This observation illustrates the computational advantage of SMT-based constraints while simultaneously highlighting a limitation of failure-driven constraint discovery: the process of encountering failures may itself be intractable.
+The motivating principle behind TAMP is that efficient planning requires the right abstraction.
+We posit that the success of VIZ-COAST stems from assigning VLMs a role aligned with their strengths: synthesizing constraints to augment the symbolic abstractions used in TAMP.
+The disparity in performance of the VLM-only baseline across domains illustrates this point.
+In Containers and Packing, this baseline achieved a 100% success rate, whereas in Blocks its success rate dropped as the number of blocks increased.
+We attribute this discrepancy to differences in what makes each domain challenging: in Containers and Packing, once key geometric constraints are recognized, planning becomes straightforward, whereas the Blocks domain still requires long-horizon, non-ergodic search.
+VLMs demonstrate strong geometric intuition but remain inefficient at this precise search.
+By isolating the VLM to the task of inferring constraints while delegating search to a symbolic planner, VIZ-COAST leverages the geometric intuition of VLMs while retaining the efficiency and precision of classical planning.
+More broadly, VIZ-COAST belongs to a growing class of approaches that leverage VLMs to make TAMP more tractable.
+These approaches suggest that learned commonsense models can complement classical planning systems by leveraging knowledge that is difficult to encode symbolically.
+In future work, it would be interesting to explore integrating constraint synthesis with complementary strategies such as VLM-TAMP, which uses VLMs to propose intermediate subgoals which reduce the search space for long-horizon tasks.
+VI-B The Strengths and Limits of VLMs in Constraint Synthesis
+Table I shows that VIZ-COAST performs robustly across a variety of VLM providers, models, and reasoning budgets.
+In general, increasing the chain-of-thought reasoning budget leads to more consistent inference of correct constraints, suggesting that deeper reasoning improves the model’s ability to extract the geometric structure required for downward refinability.
+At the same time, fast models such as Gemini-3-Flash still achieve strong performance while requiring as little as 15 seconds to generate constraints.
+VIZ-COAST aims first to act as a preventative method, proactively foreseeing and avoiding refinement failures before planning.
+However, as demonstrated in the Packing domain, VLMs are not always able to infer all necessary constraints a priori.
+Thus, in more challenging domains where initial constraints fail, VIZ-COAST leverages feedback from planning to update constraints.
+The Packing domain shows that this feedback is not merely helpful for recovering from occasional errors, but can be necessary for arriving at the correct insight.
+In particular, the no-feedback ablation was unable to solve several problem instances despite repeated attempts, whereas the feedback-enabled version succeeded.
+The ablation also encountered more refinement failures, suggesting that VIZ-COAST’s ability to reason over multiple problems produces more generalizable constraints.
+VI-C The Advantage of SMT-based Constraints
+A primary factor contributing to the high failure rates of COAST-based baselines in the Containers and Packing domains is the mechanism used to encode constraints.
+In COAST, expressing constraints requires introducing additional entities and predicates into the PDDL domain description, including timesteps and auxiliary logging and blocking predicates.
+These substantially increase the complexity of the domain representation.
+In other words, COAST’s mechanism for encoding constraints enables failure-driven constraint discovery at the expense of the utility of its abstraction.
+In contrast, VIZ-COAST encodes constraints directly through the Z3 API as additional clauses in the intermediate SMT representation.
+This allows the planner to operate on the original PDDL description while incorporating constraints with minimal additional complexity.
+In the Containers and Packing domains, this difference in representation often determined whether a planner could find a solution within seconds or fail entirely.
 VII Conclusion
-We have presented VIZ-COAST, a TAMP solver that uses VLMs’ common-sense geometric reasoning capabilities to constrain TAMP problems a priori.
-Unlike prior works, the constraints produced by the VLM generalize to unseen problem instances within a broader domain and incur minimal additional overhead by directly interfacing through APIs to an SMT-based task planner.
-These constraints eliminate potential downward refinement failures that arise between the symbolic abstraction and geometry of the world—preventing the planner from attempting to pick obstructed objects, place objects in cluttered locations, and perform other infeasible actions.
-We have demonstrated that VIZ-COAST effectively leverages VLMs’ insights to improve planning times by reducing or even eliminating replanning due to downward refinement failures.
-For example, in the Blocks domain, VIZ-COAST produced general constraints on the first attempt, resulting in task plans that were refinable on the first attempt for all 35 problem instances.
-In the Containers domain, VIZ-COAST produces constraints that enable solutions to all 30 instances, while baseline methods fail to find solutions due to the complexity of the unconstrained search space.
-Overall, VIZ-COAST demonstrates that VLMs can provide valuable insights into problem structure that accelerate search performance.
-These results are promising for future applications that extract even richer structural information, operate in online settings, or require constraint insertion based on natural language instructions grounded to the physical world.
-One limitation of our approach is that VLM inference introduces unpredictable latencies.
-Our method is agnostic to the specific VLM model used; as models become faster and more efficient, we expect this variance to decrease.
-Evaluating our methodology with local or distilled models represents an interesting direction for future work.
-We also evaluated our approach exclusively in simulation environments.
-While we believe that the generality of large pre-trained VLMs will enable constraint extraction in real-world settings, empirical validation of this capability remains future work.
+We have presented VIZ-COAST, a TAMP approach that uses VLMs’ common-sense geometric reasoning capabilities to constrain TAMP problems.
+Unlike prior works, the constraints produced generalize to unseen problem instances within a broader domain and incur minimal overhead by directly interfacing with an SMT-based task planner.
+These constraints prevent downward refinement failures when grounding the symbolic task plan to continuous motion.
+Furthermore, VIZ-COAST’s closed-loop mechanism for updating constraints with planner feedback enables reasoning over several problem instances to solve difficult domains.
+We have demonstrated that our approach improves planning times and success rates by reducing downward refinement failures and eschewing the need for auxiliary symbolic constructions required by baselines.
+For example, in the Blocks and Containers domains, VIZ-COAST produced general constraints on the first attempt, resulting in task plans that were refinable within $K$ incremental attempts for all 30 problem instances.
+In the Containers and Packing domains, VIZ-COAST solves all problems, while baseline methods often fail to find solutions due to the complexity introduced by auxiliary constructions.
+Finally, we empirically demonstrate that the performance of VIZ-COAST is robust to the VLM model used, and that current standard services are fast and powerful enough to be practical through our method.
+Overall, VIZ-COAST demonstrates that VLMs can provide valuable insights into problem structure that accelerate task-and-motion-planning.
+These results are promising for future applications that may plan using richer abstraction, operate in dynamic online settings, or require constraint insertion based on natural language instructions grounded to the physical world.
+Although we have validated the effectiveness of VLM models across several providers, our reliance on a proprietary cloud-based service remains a source of variance.
+Evaluating our method with local or distilled models is an interesting direction for future work.
+Additionally, the generalization capacity of inferred constraints is currently limited to problems within the same PDDL domain.
+Future work could explore transferring or composing constraints learned from simpler domains to solve more complex ones.
+Lastly, we evaluated our approach exclusively in simulation.
+While we believe that the generality of large pre-trained VLMs will enable constraint synthesis in a wide range of settings, empirical validation of VIZ-COAST’s ability to handle the geometric and perceptual complexity of the real world remains future work.